OpenStudy (anonymous):

The price of a Ipod dropped from $299.99 to $180.55. What was the percent decrease in price. (Rounded to nearest hundredth percent).

OpenStudy (anonymous):

decrease = 299.99 - 180.55 = 119.44 % desrease = 119.44 / 299.99 * 100 = 39.81 %